Detection system (Output Optics)Canuel, Marque - 14:29 Friday 15 May 2009 (22931)
Print this reportDetection tower interventionThis morning, we entered into the detection tower to remove all suspended baffles (4 in total). We also measured their frequency of oscillation.
Rectangular baffle 1 (beam passing through it): 0.77Hz
Rectangular baffle 2 (on the north respect to the other one; dumping B1 beams reflected by output window): 0.8Hz
Circular baffle 1 (between tower wall and rectangular baffle): 0.71Hz
Circular baffle 2 (at the jonction between cryotrap and detection tower): 1Hz